OpenStudy (anonymous):

how do i find a if my equation is y=a(x-.5)^2+2

OpenStudy (whpalmer4):

Do you have a known value of x and y?

OpenStudy (anonymous):

No. I just have my vertex for h,k which is (0.5,2)

jimthompson5910 (jim_thompson5910):

do you know any other points on this parabola?

OpenStudy (anonymous):

yes (0.5,2) & (7,2)

OpenStudy (whpalmer4):

your vertex is certainly a point on the parabola, no?

OpenStudy (anonymous):

Yes

OpenStudy (whpalmer4):

so plug in the coordinates of a known point and solve for \(a\)

OpenStudy (anonymous):

Thats where im confused. The equation in general form is y=a(x-h)^2+k my h,k can either be (7,2) or (0.5,2) so i have now y=a(x-7)^2+2 ?

jimthompson5910 (jim_thompson5910):

y=a(x-.5)^2+2 2=a(7-0.5)^2+2 ... plug in (7,2) 2=a(6.5)^2+2 ... ... ... a = ???

OpenStudy (anonymous):

A=0.04?

jimthompson5910 (jim_thompson5910):

oh wait, are you sure the other point (that's not the vertex) is (7,2) ?

OpenStudy (anonymous):

The two points on the paraobla i have is (.5, 2) & (7,2)

jimthompson5910 (jim_thompson5910):

then none of those two are the vertex

OpenStudy (anonymous):

So what is the h,k then ?

jimthompson5910 (jim_thompson5910):

the vertex would be the point (h,k) where h = (0.5+7)/2 = 7.5/2 = 3.75 but we can't determine the value of k unless we get more info

OpenStudy (anonymous):

What else information do i need

jimthompson5910 (jim_thompson5910):

another point at least

OpenStudy (anonymous):

(9,2)(2.5,2.08)(5.68,5.1)

jimthompson5910 (jim_thompson5910):

are you able to post a picture of this at all?

OpenStudy (anonymous):

How do i post it?

jimthompson5910 (jim_thompson5910):

click attach file

jimthompson5910 (jim_thompson5910):

under the text box

OpenStudy (anonymous):

Where?

jimthompson5910 (jim_thompson5910):

see where it says "draw"?

OpenStudy (anonymous):

im on my phone i dont see it

jimthompson5910 (jim_thompson5910):

oh, hmm I'm not sure if it's the same on mobile devices

jimthompson5910 (jim_thompson5910):

try scrolling down?

OpenStudy (anonymous):

I can send the link

jimthompson5910 (jim_thompson5910):

ok

OpenStudy (anonymous):

Its not letting me , but the picture is on google search real world parabolas And its the sixth picture

jimthompson5910 (jim_thompson5910):

all I see is the mcdonalds picture and a bunch of roller coasters...hmm I'm doing something wrong

OpenStudy (anonymous):

No its the tower picture after the red roller coaster

jimthompson5910 (jim_thompson5910):

so this page http://mathforum.org/mathimages/index.php/Parabola ???

OpenStudy (anonymous):

No go to google image search & type real world parabolas

jimthompson5910 (jim_thompson5910):

this is what I get http://cameronjtesblog.blogspot.com/2011/04/real-world-parabolas.html

jimthompson5910 (jim_thompson5910):

so I think I'm on the wrong page (again...)

OpenStudy (anonymous):

Go to google.com

jimthompson5910 (jim_thompson5910):

ok

jimthompson5910 (jim_thompson5910):

then what?

OpenStudy (anonymous):

Then type in real world parabolas

jimthompson5910 (jim_thompson5910):

ok did that

jimthompson5910 (jim_thompson5910):

which # link am I looking for?

OpenStudy (anonymous):

There's no link . Its a picture of a tower

jimthompson5910 (jim_thompson5910):

yeah getting the same results, are you sure there's no way to send the link? the direct link

OpenStudy (anonymous):

It wont copy right . Im trying

jimthompson5910 (jim_thompson5910):

thanks

OpenStudy (anonymous):

Do you see it

jimthompson5910 (jim_thompson5910):

I get the same results that I did if I were to search it myself I meant the direct link to the actual picture...sorry for being a pain

jimthompson5910 (jim_thompson5910):

I think your phone and my computer deliver different results or something, not sure

OpenStudy (anonymous):

Its okay hold on.

jimthompson5910 (jim_thompson5910):

ok

jimthompson5910 (jim_thompson5910):

yeah that's what I've been getting...but that has no numbers on it

jimthompson5910 (jim_thompson5910):

I'm guessing you're supposed to add in the points later?

OpenStudy (anonymous):

yeah i use this geogebra thing & it gives me the points

jimthompson5910 (jim_thompson5910):

ah send me that if you can as well, if not, then its ok

jimthompson5910 (jim_thompson5910):

basically I was hoping to see this thing for myself to see where the vertex was

OpenStudy (anonymous):

Im not close to a computer so i cant

jimthompson5910 (jim_thompson5910):

alright then no worries

OpenStudy (anonymous):

Okay

jimthompson5910 (jim_thompson5910):

I can tell you right off the bat that the vertex will not be on the same horizontal level as any other point that's on the parabola

jimthompson5910 (jim_thompson5910):

so there's no way that (0.5, 2) can be the vertex if another point is (7, 2)

OpenStudy (anonymous):

So it would be the middle point

jimthompson5910 (jim_thompson5910):

yeah at (3.75, k)

OpenStudy (anonymous):

But how do i find k

jimthompson5910 (jim_thompson5910):

and you also said that (9,2) is on the parabola right?

OpenStudy (anonymous):

Yes

jimthompson5910 (jim_thompson5910):

that's just not possible in any parabola, only 2 points will share the same y coordinate (so you can't have 3 points that have a y coordinate of 2)

jimthompson5910 (jim_thompson5910):

so this is part of the reason for my confusion and why I wanted to see a pic of the whole thing

OpenStudy (anonymous):

I have to make a quadratic formula based of the parabola in the picture & the geogebra gaves me those points

jimthompson5910 (jim_thompson5910):

hmm there's probably a typo somewhere then if we had the 3 points (0,5, 2) (7, 2) and (9,2) the you would either have a linear equation or a cubic (but not a quadratic) going through those 3 points

jimthompson5910 (jim_thompson5910):

so hopefully you see the issue here

OpenStudy (anonymous):

no the points are (9,2)(2.5,2.0)(5.68,5.1)

jimthompson5910 (jim_thompson5910):

I'm assuming those points are labeled A, B, and C in geogebra right?

OpenStudy (anonymous):

Yeah

jimthompson5910 (jim_thompson5910):

ok what you can do is with geogebra you can type "fitpoly[A,B,C,2]" without quotes into the input bar at the bottom what this will do is interpolate a polynomial (ie draw a polynomial through all 3 points) and it will force the polynomial to be of degree 2, so it will be a parabola

jimthompson5910 (jim_thompson5910):

this saves you a lot of time from doing it by hand (which involves ugly systems of equations and matrices), so that's fortunate you're able to use geogebra

OpenStudy (anonymous):

Okay. And then from there what do i do

jimthompson5910 (jim_thompson5910):

what equation do you get?

jimthompson5910 (jim_thompson5910):

when you type that command into geogebra

OpenStudy (anonymous):

I didnt do i just wanted to know so that when i do i wont get stuck

jimthompson5910 (jim_thompson5910):

well once you have the equation, you can determine the value of 'a' by looking at the leading coefficient

jimthompson5910 (jim_thompson5910):

you can also determine the vertex by typing in "max[f,-20,20]" without quotes into the input bar

OpenStudy (anonymous):

Okay, & then.from there im asked to find the x & y-intercepts

jimthompson5910 (jim_thompson5910):

the y-intercept would be the last number in the function you get from after doing the "fitpoly" command the x-intercepts can be found by typing "root[f]" without quotes into the input bar

OpenStudy (anonymous):

Okay thanks

jimthompson5910 (jim_thompson5910):

you're welcome